# Probabilistic tube-based control synthesis of stochastic multi-agent systems under signal temporal logic

### v1

@@ I Introduction
Multi-agent systems (MASs) can be found in many applications, such as robotics, autonomous vehicles, and cyber-physical systems. When these systems are stochastic, the formal specification of system properties can be formulated in a probabilistic setting, enabling a trade-off between quantifying uncertainty and adjusting feasibility. As the complexity in control synthesis from temporal logic under uncertainty grows with the dimensionality of the overall system, existing approaches typically focus on single-agent schemes [1] or non-stochastic systems [2, 3].
@@
In this paper, we focus on signal temporal logic (STL) [4], as a tool to formally formulate and verify specifications for a wide range of MAS applications. STL employs predicates as atomic elements, coupled with Boolean and temporal operators, allowing precise specification of complex spatio-temporal properties in a dynamical system. In a deterministic setting, it is possible to design sound and complete control algorithms that guarantee STL satisfaction [5], based on the quantitative semantics of STL [6]. Here we consider stochastic MASs and a stochastic optimal control problem, where the goal is to satisfy a multi-agent STL specification with a predefined probability.
@@
To address stochasticity in the STL framework, the works in [7, 8, 9] propose risk constraints over predicates while preserving Boolean and temporal operators. Probabilistic STL in [10] allows one to express uncertainty by incorporating random variables into predicates, while [11] introduces chance-constrained temporal logic for modeling uncertainty in autonomous vehicles. Similar approaches are found in [12, 13]. Top-down approaches imposing chance constraints on the entire specification are explored in [1, 14, 15]. Although important, these works focus on single-agent or low-dimensional systems and lack guidance on extending to MASs. An extension of [1] to stochastic MASs under STL has recently been presented in [16], considering, however, only a single joint task per agent and bounded distributions.
@@
Here, we propose a two-stage approach to solve a stochastic optimal control problem for discrete-time linear MASs subject to additive stochastic perturbations and a global STL specification permitting multiple individual and joint tasks per agent. First, we decompose the multi-agent dynamics into a deterministic system and an error closed-loop stochastic system, for which we construct a probabilistic reachable tube (PRT) [17] as the Cartesian product of reachable sets of individual error systems. These are driven by stochastic disturbances within confidence regions (CRs) with a fixed probability. By assuming independence among individual disturbances, we show that the PRT probability can be controlled by the product of probabilities selected for each individual CR and a union-bound argument applied over time. Thus, by lower bounding the PRT probability by the specification probability, we can tighten all state constraints induced by the STL specification by solving tractable optimization problems over segments of the PRT. For multi-agent STL specifications, this is a less conservative alternative to tightening approaches relying on the STL structure [1, 16]. An attainable feasible solution to the resulting deterministic problem can then be used to synthesize multi-agent trajectories that satisfy the STL specification with the desired confidence level. To the best of the authors’ knowledge, this work is the first to address stochastic MASs under STL utilizing PRTs. Subsequently, to enhance scalability, we decompose the resulting deterministic problem into agent-level subproblems, which are solved iteratively according to a scheduling policy. We show that this iterative procedure is recursively feasible, ensures satisfaction of local tasks, and guarantees nondecreasing robustness for joint tasks.
@@ II-A Notation and Preliminaries
Based on the semantics introduced above, the horizon of a formula is recursively defined as [4]: $N^{\pi}=0$, $N^{\lnot\phi}=N^{\phi}$, $N^{\phi_{1}\land\phi_{2}}=\max(N^{\phi_{1}},N^{\phi_{2}})$, $N^{\phi_{1}\;U_{[a,b]}\phi_{2}}=b+\max(N^{\phi_{1}},N^{\phi_{2}})$.
STL is endowed with robustness metrics for assessing the robust satisfaction of a formula [6]. A scalar-valued function $\rho^{\phi}:{\rm I\!R}^{n}\times\cdots\times{\rm I\!R}^{n}\to{\rm I\!R}$ of a signal indicates how robustly a signal $\bm{x}(t)$ satisfies a formula $\phi$. The robustness function is defined recursively as follows: $\rho^{\pi}(\bm{x}(t))=\mu(x(t))$, $\rho^{\lnot\phi}(\bm{x}(t))=-\rho^{\phi}(\bm{x}(t))$,
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph containing no self-loops, with node set $\mathcal{V}$, cardinality $M=|\mathcal{V}|$, and edge set $\mathcal{E}$. Let also $\mathcal{V}^{\prime}\subseteq\mathcal{V}$, with $|\mathcal{V}^{\prime}|>1$, and define $\mathcal{E}_{\mathcal{V}^{\prime}}\subseteq\mathcal{E}$ as the set of edges attached to nodes $\mathcal{V}^{\prime}$. Then, $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}_{\mathcal{V}^{\prime}})$ is a clique [18], i.e., a complete subgraph of $\mathcal{G}$, if $\mathcal{E}_{\mathcal{V}^{\prime}}$ contains all possible edges between nodes $\mathcal{V}^{\prime}$. The set of cliques of $\mathcal{G}$ is defined as
@@ III-A Error dynamics and construction of probabilistic tubes
Next, we show how to obtain $t$-PRSs for (8b) at probability levels $\hat{\theta}_{t}$, $t\in{\rm I\!N}_{[0,N]}$, by identifying subsets of $\mathcal{W}$ with a degree of concentration of $\mathcal{D}_{w}$ [17].
@@
Let now $\mathcal{D}_{e_{i}(t)}$ be the distribution of $e_{i}(t)$, and $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t)})$ be a CR for $e_{i}(t)$ at probability $\theta_{i}$. Since, $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(0)})\subseteq E_{i}(0)=\{0\}$, we have $\bar{A}_{i}\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e(0)})\oplus\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})\subseteq\bar{A}_{i}E_{i}(0)\oplus\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})=E_{i}(1)$, so $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(1)})\subseteq E_{i}(1)$, as $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t+1)})\subseteq\bar{A}_{i}\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t)})\oplus\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})$ for all $t\in{\rm I\!N}$, since $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})$ is an ellipsoidal region by [17, Cor. 4]. Inductively we show that $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t)})\subseteq E_{i}(t)$, $i\in{\rm I\!N}_{[1,M]}$, for all $t\in{\rm I\!N}$. The latter implies that $\mathrm{Pr}\{e_{i}(t)\in E_{i}(t)\}\geq\theta_{i}$, $i\in{\rm I\!N}_{[1,M]}$, from which we have $\mathrm{Pr}\{e(t)\in E(t)\}=\mathrm{Pr}\{(e_{1}(t)\in E_{1}(t))\wedge\cdots\wedge(e_{M}(t)\in E_{M}(t))=\mathrm{Pr}\{(e_{1}(t)\in E_{1}(t))\}\cdots\mathrm{Pr}\{(e_{M}(t)\in E_{M}(t))\}\geq\theta_{1}\theta_{2}\cdots\theta_{M}$, which follows from the independence of $E_{i}(t)$, $i\in{\rm I\!N}_{[1,M]}$, by construction, for all $t\in{\rm I\!N}$. ∎
@@ III-B Constraint tightening
Clearly, apart from the volume of $\bm{E}$, the gain $K$ affects the feasible domain of (11) too. Due to space limitations, we will present its construction in future work. By selecting $\|{}\cdot{}\|_{1}$-based costs, problem (11) can be formulated as a mixed-integer linear program (MILP) using binary encoding for STL constraints [5]. The complexity of (11), which grows with the dimension of the MAS, is addressed in the next section.
@@ III-C2 Iterative algorithm
at $k=0$ for $i\in{\rm I\!N}_{[1,M]}$. Note that by Assum. 3, Problem (14) is guaranteed to be feasible. After performing the above initialization routine, at iteration $k\geq 1$, only a subset of agents, denoted by $\mathcal{O}_{k}{\subset}\mathcal{V}$, are allowed to update their input sequences by performing an optimization, while the remaining agents retrieve their input sequences from the previous iteration $k-1\geq 0$. Roughly, the set $\mathcal{O}_{k}\subset\mathcal{V}$ is constructed so that any combination of its elements does not belong to a clique $\nu\in\mathcal{K}_{\phi}$. Due to space limitations, we simply construct $\mathcal{O}_{k}$ as a singleton, which only affects the number of agents’ trajectories that can be optimized in parallel per iteration. For the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $\mathcal{V}={\rm I\!N}_{[1,M]}$, $\mathcal{O}_{k}=\mathrm{mod}(\mathcal{O}_{k-1},M)+1$, for $k>1$, with $\mathcal{O}_{1}=1$. We refer readers to [19, Sec. V.B] for a more efficient construction of $\mathcal{O}_{k}$ enabling parallel computations at each iteration (see Sec. IV for a numerical example).
@@ IV Example
After we have derived $\psi$, we formulate (11) as an MILP using the YALMIP toolbox [20], in MATLAB. However, possibly due to the high dimensionality of the problem, the GUROBI solver [21] could not return a solution in a reasonable time. As a remedy, we implement the iterative optimization procedure presented in Section III-C, after decomposing $\psi$ according to (13), based on the sets $\mathcal{T}_{1}=\{(2,3),5\}$, $\mathcal{T}_{2}=\{(1,3)\}$, $\mathcal{T}_{3}=\{(1,2),4\}$, $\mathcal{T}_{4}=\{3,5\}$, $\mathcal{T}_{5}=\{1,4,6\}$, $\mathcal{T}_{6}=\{5,8,9\}$, $\mathcal{T}_{7}=\{4,8\}$, $\mathcal{T}_{8}=\{6,7,10\}$, $\mathcal{T}_{9}=\{6,10\}$, and $\mathcal{T}_{10}=\{8,9\}$. By selecting sets $\mathcal{O}_{k}$, $k>1$, as $\mathcal{O}_{1}=\{1,4,6,10\}$, $\mathcal{O}_{2}=\{3,5,7,9\}$, $\mathcal{O}_{3}=\{8,9,2,5\}$, $\mathcal{O}_{4}=\mathcal{O}_{1}$, $\mathcal{O}_{5}=\mathcal{O}_{2}$, $\mathcal{O}_{6}=\mathcal{O}_{3}$, $\mathcal{O}_{7}=\mathcal{O}_{1}$, and so on, we run Algorithm 1, which terminates after seven iterations returning a multi-agent trajectory that satisfies the global STL task $\psi$. Computational aspects of Algorithm 1 are depicted in Fig. 3 (left), while Fig. 3 (right) illustrates the evolution of the robustness function computed at the agent level during the execution of Algorithm 1. Fig. 4 displays $10^{3}$ realizations of noisy multi-agent trajectories, illustrating also the impact of constraint tightening on the workspace $\mathcal{X}$, obstacles $O_{1}$, $O_{2}$, $O_{3}$, and the regions $T_{i}$, $G_{i}$ ($i\in{\rm I\!N}_{[1,10]}$). By evaluating the robustness function of the global STL specification $\phi$ for numerous noisy realizations, we see that $\phi$ is violated in less than $30\%$ of the total realizations, verifying Theorem 2.

### v2

@@ I Introduction
Multi-agent systems (MASs) can be found in many applications, such as robotics, autonomous vehicles, and cyber-physical systems. When these systems are stochastic, the formal specification of system properties can be formulated in a probabilistic setting, enabling a trade-off between quantifying uncertainty and adjusting feasibility. As the complexity in control synthesis from temporal logic under uncertainty grows with the dimensionality of the overall system, existing approaches typically focus on single-agent schemes [FarahaniTAC2019] or non-stochastic systems [Liu2017, Buyukkocak2021].
@@
In this paper, we focus on signal temporal logic (STL) [MalerSTL2004], as a tool to formally formulate and verify specifications for a wide range of MAS applications. STL employs predicates as atomic elements, coupled with Boolean and temporal operators, allowing precise specification of complex spatio-temporal properties in a dynamical system. In a deterministic setting, it is possible to design sound and complete control algorithms that guarantee STL satisfaction [MurrayCDC2014], based on the quantitative semantics of STL [Donze2010]. Here we consider stochastic MASs and a stochastic optimal control problem, where the goal is to satisfy a multi-agent STL specification with a predefined probability.
@@
To address stochasticity in the STL framework, the works in [Safaoui2020, LindemannCDC2020, LindemannTAC2022] propose risk constraints over predicates while preserving Boolean and temporal operators. Probabilistic STL in [Sadigh2016] allows one to express uncertainty by incorporating random variables into predicates, while [Sadigh2018] introduces chance-constrained temporal logic for modeling uncertainty in autonomous vehicles. Similar approaches are found in [Li2017, Kyriakis2019]. Top-down approaches imposing chance constraints on the entire specification are explored in [FarahaniTAC2019, ScherHSCC2022, Scher2022]. Although important, these works focus on single-agent or low-dimensional systems and lack guidance on extending to MASs. An extension of [FarahaniTAC2019] to stochastic MASs under STL has recently been presented in [Yang2023], considering, however, only a single joint task per agent and bounded distributions.
@@
Here, we propose a two-stage approach to solve a stochastic optimal control problem for discrete-time linear MASs subject to additive stochastic perturbations and a global STL specification permitting multiple individual and joint tasks per agent. First, we decompose the multi-agent dynamics into a deterministic system and an error closed-loop stochastic system, for which we construct a probabilistic reachable tube (PRT) [HewingECC2018] as the Cartesian product of reachable sets of individual error systems. These are driven by stochastic disturbances within confidence regions (CRs) with a fixed probability. By assuming independence among individual disturbances, we show that the PRT probability can be controlled by the product of probabilities selected for each individual CR and a union-bound argument applied over time. Thus, by lower bounding the PRT probability by the specification probability, we can tighten all state constraints induced by the STL specification by solving tractable optimization problems over segments of the PRT. For multi-agent STL specifications, this is a less conservative alternative to tightening approaches relying on the STL structure [FarahaniTAC2019, Yang2023]. An attainable feasible solution to the resulting deterministic problem can then be used to synthesize multi-agent trajectories that satisfy the STL specification with the desired confidence level. To the best of the authors’ knowledge, this work is the first to address stochastic MASs under STL utilizing PRTs. Subsequently, to enhance scalability, we decompose the resulting deterministic problem into agent-level subproblems, which are solved iteratively according to a scheduling policy. We show that this iterative procedure is recursively feasible, ensures satisfaction of local tasks, and guarantees nondecreasing robustness for joint tasks.
@@ II-A Notation and Preliminaries
Based on the semantics introduced above, the horizon of a formula is recursively defined as [MalerSTL2004]: $N^{\pi}=0$, $N^{\lnot\phi}=N^{\phi}$, $N^{\phi_{1}\land\phi_{2}}=\max(N^{\phi_{1}},N^{\phi_{2}})$, $N^{\phi_{1}\;U_{[a,b]}\phi_{2}}=b+\max(N^{\phi_{1}},N^{\phi_{2}})$.
STL is endowed with robustness metrics for assessing the robust satisfaction of a formula [Donze2010]. A scalar-valued function $\rho^{\phi}:{\rm I\!R}^{n}\times\cdots\times{\rm I\!R}^{n}\to{\rm I\!R}$ of a signal indicates how robustly a signal $\bm{x}(t)$ satisfies a formula $\phi$. The robustness function is defined recursively as follows: $\rho^{\pi}(\bm{x}(t))=\mu(x(t))$, $\rho^{\lnot\phi}(\bm{x}(t))=-\rho^{\phi}(\bm{x}(t))$,
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph containing no self-loops, with node set $\mathcal{V}$, cardinality $M=|\mathcal{V}|$, and edge set $\mathcal{E}$. Let also $\mathcal{V}^{\prime}\subseteq\mathcal{V}$, with $|\mathcal{V}^{\prime}|>1$, and define $\mathcal{E}_{\mathcal{V}^{\prime}}\subseteq\mathcal{E}$ as the set of edges attached to nodes $\mathcal{V}^{\prime}$. Then, $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}_{\mathcal{V}^{\prime}})$ is a clique [Orlin1977], i.e., a complete subgraph of $\mathcal{G}$, if $\mathcal{E}_{\mathcal{V}^{\prime}}$ contains all possible edges between nodes $\mathcal{V}^{\prime}$. The set of cliques of $\mathcal{G}$ is defined as
@@ III-A Error dynamics and construction of probabilistic tubes
Next, we show how to obtain $t$-PRSs for (8b) at probability levels $\hat{\theta}_{t}$, $t\in{\rm I\!N}_{[0,N]}$, by identifying subsets of $\mathcal{W}$ with a degree of concentration of $\mathcal{D}_{w}$ [HewingECC2018].
@@
Let now $\mathcal{D}_{e_{i}(t)}$ be the distribution of $e_{i}(t)$, and $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t)})$ be a CR for $e_{i}(t)$ at probability $\theta_{i}$. Since, $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(0)})\subseteq E_{i}(0)=\{0\}$, we have $\bar{A}_{i}\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e(0)})\oplus\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})\subseteq\bar{A}_{i}E_{i}(0)\oplus\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})=E_{i}(1)$, so $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(1)})\subseteq E_{i}(1)$, as $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t+1)})\subseteq\bar{A}_{i}\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t)})\oplus\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})$ for all $t\in{\rm I\!N}$, since $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{w_{i}})$ is an ellipsoidal region by [HewingECC2018, Cor. 4]. Inductively we show that $\mathscr{E}_{\theta_{i}}(\mathcal{D}_{e_{i}(t)})\subseteq E_{i}(t)$, $i\in{\rm I\!N}_{[1,M]}$, for all $t\in{\rm I\!N}$. The latter implies that $\mathrm{Pr}\{e_{i}(t)\in E_{i}(t)\}\geq\theta_{i}$, $i\in{\rm I\!N}_{[1,M]}$, from which we have $\mathrm{Pr}\{e(t)\in E(t)\}=\mathrm{Pr}\{(e_{1}(t)\in E_{1}(t))\wedge\cdots\wedge(e_{M}(t)\in E_{M}(t))=\mathrm{Pr}\{(e_{1}(t)\in E_{1}(t))\}\cdots\mathrm{Pr}\{(e_{M}(t)\in E_{M}(t))\}\geq\theta_{1}\theta_{2}\cdots\theta_{M}$, which follows from the independence of $E_{i}(t)$, $i\in{\rm I\!N}_{[1,M]}$, by construction, for all $t\in{\rm I\!N}$. ∎
@@ III-B Constraint tightening
Clearly, apart from the volume of $\bm{E}$, the gain $K$ affects the feasible domain of (11) too. Due to space limitations, we will present its construction in future work. By selecting $\|{}\cdot{}\|_{1}$-based costs, problem (11) can be formulated as a mixed-integer linear program (MILP) using binary encoding for STL constraints [MurrayCDC2014]. The complexity of (11), which grows with the dimension of the MAS, is addressed in the next section.
@@ III-C2 Iterative algorithm
at $k=0$ for $i\in{\rm I\!N}_{[1,M]}$. Note that by Assum. 3, Problem (14) is guaranteed to be feasible. After performing the above initialization routine, at iteration $k\geq 1$, only a subset of agents, denoted by $\mathcal{O}_{k}{\subset}\mathcal{V}$, are allowed to update their input sequences by performing an optimization, while the remaining agents retrieve their input sequences from the previous iteration $k-1\geq 0$. Roughly, the set $\mathcal{O}_{k}\subset\mathcal{V}$ is constructed so that any combination of its elements does not belong to a clique $\nu\in\mathcal{K}_{\phi}$. Due to space limitations, we simply construct $\mathcal{O}_{k}$ as a singleton, which only affects the number of agents’ trajectories that can be optimized in parallel per iteration. For the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $\mathcal{V}={\rm I\!N}_{[1,M]}$, $\mathcal{O}_{k}=\mathrm{mod}(\mathcal{O}_{k-1},M)+1$, for $k>1$, with $\mathcal{O}_{1}=1$. We refer readers to [vlahakisECC2023, Sec. V.B] for a more efficient construction of $\mathcal{O}_{k}$ enabling parallel computations at each iteration (see Sec. IV for a numerical example).
@@ IV Example
After we have derived $\psi$, we formulate (11) as an MILP using the YALMIP toolbox [Yalmip], in MATLAB. However, possibly due to the high dimensionality of the problem, the GUROBI solver [gurobi] could not return a solution in a reasonable time. As a remedy, we implement the iterative optimization procedure presented in Section III-C, after decomposing $\psi$ according to (13), based on the sets $\mathcal{T}_{1}=\{(2,3),5\}$, $\mathcal{T}_{2}=\{(1,3)\}$, $\mathcal{T}_{3}=\{(1,2),4\}$, $\mathcal{T}_{4}=\{3,5\}$, $\mathcal{T}_{5}=\{1,4,6\}$, $\mathcal{T}_{6}=\{5,8,9\}$, $\mathcal{T}_{7}=\{4,8\}$, $\mathcal{T}_{8}=\{6,7,10\}$, $\mathcal{T}_{9}=\{6,10\}$, and $\mathcal{T}_{10}=\{8,9\}$. By selecting sets $\mathcal{O}_{k}$, $k>1$, as $\mathcal{O}_{1}=\{1,4,6,10\}$, $\mathcal{O}_{2}=\{3,5,7,9\}$, $\mathcal{O}_{3}=\{8,9,2,5\}$, $\mathcal{O}_{4}=\mathcal{O}_{1}$, $\mathcal{O}_{5}=\mathcal{O}_{2}$, $\mathcal{O}_{6}=\mathcal{O}_{3}$, $\mathcal{O}_{7}=\mathcal{O}_{1}$, and so on, we run Algorithm 1, which terminates after seven iterations returning a multi-agent trajectory that satisfies the global STL task $\psi$. Computational aspects of Algorithm 1 are depicted in Fig. 3 (left), while Fig. 3 (right) illustrates the evolution of the robustness function computed at the agent level during the execution of Algorithm 1. Fig. 4 displays $10^{3}$ realizations of noisy multi-agent trajectories, illustrating also the impact of constraint tightening on the workspace $\mathcal{X}$, obstacles $O_{1}$, $O_{2}$, $O_{3}$, and the regions $T_{i}$, $G_{i}$ ($i\in{\rm I\!N}_{[1,10]}$). By evaluating the robustness function of the global STL specification $\phi$ for numerous noisy realizations, we see that $\phi$ is violated in less than $30\%$ of the total realizations, verifying Theorem 2.